# Evaluating NeRFs for 3D Plant Geometry Reconstruction in Field Conditions

### v2

@@ 2.4.1. Pipline
We reconstruct the scene and capture point clouds using a FARO scan for ground truth. To perform a one-to-one comparison between the NeRF-based reconstruction and ground truth, alignment is crucial. Our alignment and evaluation methodology is adapted from Knapitsch et al. [Knapitsch et al., 2017]. In their work, they evaluate different pipelines and use colmmap as a ’arbitrary reference’ frame. However, in our case, all the NeRFs use colmap in their pipeline so the reference frame and reconstruction frame becomes the same. This simplifies the steps and the updated steps, used in our paper are given below:
  1. (1)
@@
  2. (2)
    Cropping: Each ground-truth model has a manually-defined bounding volume, outlining the evaluation region for reconstructions.
  3. (3)
    ICP Registration: Drawing inspiration from the iterative refinement process detailed by Besl et al. [Besl and McKay, 1992] and further refined by Zhang et al. [Zhang, 1994], we adopt a three-stage approach as introduced by Knapitsch et al. [Knapitsch et al., 2017] for our initial registration framework. The process begins with a specified voxel size and an associated threshold for the initial registration. In the next iteration, the transformation result from the previous step is used as a starting point, with the voxel size reduced by half to achieve finer detail in the registration. The third stage aims to refine the alignment further by returning to the original voxel size and adjusting the threshold to facilitate convergence at each stage. This multi-scale strategy is designed to capture both coarse and fine details, thereby improving the accuracy and precision of the model alignment. However, in our adaptation for plant structure reconstruction, we diverged from Knapitsch et al. by maintaining the iterative process within a single stage rather than expanding across multiple stages. We found that increasing the iteration count tenfold, rather than the number of stages, prevented the registration process from collapsing [Billings et al., 2015].
2.4.2. Metrics
@@
    $P(d)=\frac{100}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}\mathbb{I}\left(\min_{\mathbf{g}\in\mathcal{G}}\|\mathbf{r}-\mathbf{g}\|<d\right),$
    where $\mathbb{I}(\cdot)$ is an indicator function.
  2. (2)
@@
    $R(d)=\frac{100}{|\mathcal{G}|}\sum_{\mathbf{g}\in\mathcal{G}}\mathbb{I}\left(\min_{\mathbf{r}\in\mathcal{R}}\|\mathbf{g}-\mathbf{r}\|<d\right).$
    Both the above two metrics are extensively utilized in recent studies. [Balloni et al., 2023, Mazzacca et al., 2023].
  3. (3)
@@
    $F(d)=\frac{2\times P(d)\times R(d)}{P(d)+R(d)}.$
@@
    The harmonic nature of the F-score ensures that if either $P(d)$ or $R(d)$ approaches zero, the F-score will also tend towards zero, providing a more robust summary statistic than the arithmetic mean. The value of $d$ was considered 0.005 for first two scenarios and 0.01 for the last scenario.
  4. (4)
    Traditional metrics such as, SSIM [Wang et al., 2004], PSNR [Hore and Ziou, 2010], LPIPS [Zhang et al., 2018] are used to evaluate the quality of the rendered images. These metrics do not need the 3D ground truth and are widely used in literature [Xu et al., 2022, Zhang et al., 2021] for evaluation.
(a) Ground Truth (LiDAR)
(b) Reconstruction (NeRF)
(c) Precision
(d) Recall
Figure 5. Visualization for Metrics
2.4.3. Visualization Color Interpretation
The color-coded visualizations employed in our research provide an intuitive understanding of spatial relationships within the 3D reconstructed plant structures. The color gradients signify varying distances from a specific reference point, enabling a comprehensive assessment of depth and intricacy. Specifically:
  1. (1)
    Grey: Represents regions of the plant structure that are closer or have the least distance to a particular reference.
  2. (2)
    Red: Depicts intermediate distances of points from the reference pointcloud within the plant structure.
  3. (3)
    Black: Highlights the furthest points or regions within the reconstructed structure.
3. Results
The results from all the scenarios are provided below:
3.1. Scenario I - Single Plants Indoors
Figure 6. Camera Pose Estimations for Scenario-I
The estimated camera poses from colmap are visualized in Fig. 6. The results from our evaluation of three state-of-the-art NeRF techniques for 3D plant structure reconstruction are illustrated in the Fig. 7.
@@ 3.1. Scenario I - Single Plants Indoors
(d) Precision
(e) Recall
(f) F1 Score
Number of Iterations
  Instant-NGP    NeRFacto    TensoRF
Figure 7. Combined Comparison of 2D Image Quality (top) and 3D Geometry (bottom) Metrics for Scenario-I
Precision:
Across all models, precision generally increases with the number of iterations. Instant-NGP shows a significant leap in precision from 100 to 5000 iterations (0.29 to 21.93), indicating a drastic improvement in the accuracy of reconstructed points relative to the ground truth. NeRFacto demonstrates a more consistent and steep rise in precision, reaching a peak of 73.57 at 30000 iterations, which surpasses Instant-NGP’s best precision. TensoRF, however, shows a relatively modest increase in precision, suggesting its limited capability in accurately capturing fine details compared to the other two models. It is further illustrated in Fig. 8.
@@
Figure 8. Precision of 3D reconstruction of NeRF techniques
Recall:
The recall metric follows a similar trend, with Instant-NGP and NeRFacto showing substantial increases with more iterations, indicating an enhanced ability to encompass points from the ground truth. Notably, NeRFacto achieves remarkably high recall values (over 90) at higher iterations (10000 and above), suggesting its superiority in completeness of reconstruction. TensoRF’s recall values are significantly lower, suggesting that it may miss more details from the ground truth compared to the other models. It is further illustrated in Fig. 9.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 9. Recall of 3D reconstruction of NeRF techniques
F1 Score:
The F1 score, balancing precision and recall, highlights NeRFacto as the most balanced model, especially at higher iterations, with scores above 80. Instant-NGP shows a significant improvement in F1 scores as iterations increase, but it doesn’t reach the heights of NeRFacto. TensoRF lags behind in this metric, indicating a less balanced performance between precision and recall.
PSNR:
The Peak Signal-to-Noise Ratio (PSNR) reflects the quality of rendered images. In this metric, Instant-NGP and NeRFacto show a gradual increase in PSNR with more iterations, suggesting improved image quality. TensoRF’s PSNR values are lower, indicating potentially lower image quality throughout its iterations.
SSIM:
The Structural Similarity Index (SSIM) is another measure of image quality, assessing the perceived change in structural information. Here, NeRFacto and Instant-NGP both show a steady increase in SSIM with more iterations, with NeRFacto achieving slightly higher scores, suggesting better preservation of structural information in its renderings. TensoRF, again, shows relatively lower SSIM scores.
LPIPS:
The Lower Perceptual Image Patch Similarity (LPIPS) metric indicates perceived image similarity, with lower values being better. NeRFacto and Instant-NGP both show a significant decrease in LPIPS with more iterations, indicating improved perceptual similarity to the ground truth. TensoRF’s LPIPS values are consistently higher, suggesting lower perceptual similarity.
Computation Time:
Time efficiency is a crucial factor, especially for practical applications. Instant-NGP demonstrates a relatively balanced approach between efficiency and performance, with time increments correlating reasonably with the increase in iterations. However, it becomes significantly time-consuming at very high iterations (20000 and 30000). NeRFacto, while showing superior performance in many metrics, demands considerably more time, especially at higher iterations, which could be a limiting factor in time-sensitive scenarios. TensoRF, despite its lower performance in other metrics, maintains a more consistent time efficiency, suggesting its suitability for applications where time is a critical constraint. For the best performing model, it is further illustrated in Fig. 10
(a) 200 Iterations
(b) 400 Iterations
(c) 800 Iterations
(d) 5000 Iterations
(e) 10000 Iterations
(f) 30000 Iterations
Figure 10. Precision Over Training Iterations for NeRFacto - Scenario 1
Overall Performance and Suitability:
In sum, NeRFacto emerges as the most robust model in terms of precision, recall, F1 score, and image quality metrics (PSNR, SSIM, LPIPS), making it highly suitable for applications demanding high accuracy and completeness in 3D modeling. However, its time inefficiency at higher iterations might restrict its use in time-sensitive contexts. Instant-NGP presents a good balance between performance and efficiency, making it a viable option for moderately demanding scenarios. Detailed results are given in Table 3, after complete training. For more granular look, consult the supplementary Table 6. The Precision-Recall Curves based on varying distance threshold after maximum training of 30,000 iterations is given in Fig. 11
Table 3. Performance metrics of NeRFs reconstruction techniques - Scenario I
Model Name Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 24.66 90.62 38.77 23.41 0.81 0.17 756
NeRFacto 73.57 94.72 82.81 22.24 0.73 0.12 1938
TensoRF 9.58 43.34 15.69 14.69 0.55 0.66 1973
# of points (%)
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Meters
@@
Figure 11. Precision-Recall Curves (Scenario-I) based on varying distance threshold after 30,000 iterations
Insight 1: Computational Cost and Accuracy Trade-off in Instant-NGP and NeRFacto:
@@
The steep increase in performance metrics with the number of iterations for both Instant-NGP and NeRFacto suggests that these models require a substantial amount of data processing to achieve high accuracy, which is critical in high-fidelity 3D modeling. However, this also implies a higher computational cost, which needs to be considered in practical applications.
Insight 2: Model Suitability in High-Detail 3D Reconstructions:
@@
The significant disparity in the performance of TensoRF compared to the other two models, particularly in precision and recall, indicates that not all NeRF models are equally suited for tasks requiring high-detail 3D reconstructions. This highlights the importance of model selection based on the specific requirements of the application.
Insight 3: Divergence in 2D Image Quality and 3D Reconstruction in Instant-NGP:
@@
A detailed examination reveals that Instant-NGP demonstrates notable strength in 2D image quality metrics such as PSNR, SSIM, and LPIPS, reflecting its ability to produce superior rendered image quality. However, this excellence in 2D imaging does not correspondingly extend to 3D reconstruction metrics like Precision, Recall, and F1 Score. This observation highlights a significant distinction in the challenges associated with optimizing for high-quality image rendering as opposed to achieving accurate 3D representations. The model’s adeptness at rendering highly detailed 2D images does not necessarily imply its effectiveness in accurately reconstructing complex 3D structures, particularly in the context of intricate plant models. This insight underscores the need for a nuanced approach in evaluating the performance of models that are tasked with both 2D image rendering and 3D spatial reconstruction.
3.2. Scenario II - Multiple Plants Indoors

### v3

@@ 2.2   Neural Radiance Fields (NeRFs)
We reconstruct the scene and capture point clouds using a FARO scan for ground truth. 3D registration or alignment is crucial to perform a one-to-one comparison between the NeRF-based reconstruction and ground truth. Our alignment and evaluation methodology is adapted from Knapitsch et al. [44]. In their work, they evaluate different pipelines and use COLMAP as a ’arbitrary reference’ frame. However, in our case, all the NeRFs use COLMAP in their pipeline, so the reference and reconstruction frames become the same. The steps used for registration are:
@@
  • Cropping: Each ground-truth model has a manually-defined bounding volume, outlining the evaluation region for reconstruction.
  • Iterative Closest Point (ICP) Registration: Drawing inspiration from the iterative refinement process detailed by Besl and McKay [45] and further refined by Zhang [46], we adopt a three-stage approach [44] for our initial registration framework. The process begins with a specified voxel size and an associated threshold for the initial registration. In the next iteration, the transformation result from the previous step is used as a starting point, with the voxel size reduced by half to achieve finer detail in the registration. The third stage aims to refine the alignment further by returning to the original voxel size and adjusting the threshold to facilitate convergence at each stage. This multi-scale strategy is designed to capture both coarse and fine details, thereby improving the accuracy and precision of the model alignment. However, in our adaptation for plant structure reconstruction, we diverged from Knapitsch et al. [44] by maintaining the iterative process within a single stage rather than expanding across multiple stages. We found that increasing the iteration count tenfold, rather than the number of stages, prevented the registration process from collapsing [47].
2.4   Evaluation Metrics
@@
    $P(d)=\frac{100}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}\mathbb{I}\left(\min_{\mathbf{g}\in\mathcal{G}}\|\mathbf{r}-\mathbf{g}\|<d\right),$ (3)
    where $\mathbb{I}(\cdot)$ is an indicator function. Precision values ranges from 0 to 100, with higher values indicating better performance.
  2. 2.
@@
    $R(d)=\frac{100}{|\mathcal{G}|}\sum_{\mathbf{g}\in\mathcal{G}}\mathbb{I}\left(\min_{\mathbf{r}\in\mathcal{R}}\|\mathbf{g}-\mathbf{r}\|<d\right).$ (4)
    Its value ranges from 0 to 100, with higher values indicating better performance. Both the above two metrics are extensively utilized in recent studies. [35, 48].
  3. 3.
@@
    $F(d)=\frac{2\times P(d)\times R(d)}{P(d)+R(d)}.$ (5)
@@
    The harmonic nature of the F-score ensures that if either $P(d)$ or $R(d)$ approaches zero, the F-score will also tend towards zero, providing a more robust summary statistic than the arithmetic mean. F-score ranges from 0 to 100, with higher values indicating better performance. The details about value of $d$ cutoff is given later in discussion about precision-recall curves.
    For quantifying the quality of the NeRF-rendered 2D image compared to the validation image (left out from NeRF training), the following metrics are used:
  4. 4.
    Learned Perceptual Image Patch Similarity (LPIPS) [49]: To quantify the perceptual differences between two image patches, $x$ and $x_{0}$, the Learned Perceptual Image Patch Similarity (LPIPS) framework employs activations from a neural network $F$. Features are extracted from $L$ layers and normalized across the channel dimension. For each layer $l$, the normalized features are represented by $\hat{y}^{l}$ and $\hat{y}^{l0}$, which exist in the space $\mathbb{R}^{H_{l}\times W_{l}\times C_{l}}$. These are then weighted channel-wise by a vector $w_{l}\in\mathbb{R}^{C_{l}}$. The perceptual distance is computed using the $\ell_{2}$ norm, both spatially and across channels, as expressed in the equation:
    $d(x,x_{0})=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\left\|w_{l}\odot(\hat{y}^{l}_{hw}-\hat{y}^{l0}_{hw})\right\|_{2}^{2}$ (6)
    This distance metric, $d(x,x_{0})$, provides a scalar value indicating the perceptual dissimilarity between the patches. The vector $w_{l}$ weights the contribution of each channel to the distance metric. By setting $w_{l}$ to $1/\sqrt{C_{l}}$, the computation effectively measures the cosine distance, highlighting the directional alignment of the feature vectors instead of their magnitude. Its value ranges from 0 to 1, with lower values indicating better performance.
  5. 5.
    Peak Signal-to-Noise Ratio (PSNR): [50] The PSNR between two images, one being the reference and the other the reconstructed image, is defined as:
    $\text{PSNR}=10\cdot\log_{10}\left(\frac{\text{MAX}_{I}^{2}}{\text{MSE}}\right),$ (7)
    where $\text{MAX}_{I}$ is the maximum possible pixel value of the image, and MSE is the Mean Squared Error between the reference and the reconstructed image. The MSE is given by:
    $\text{MSE}=\frac{1}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\left(I(i,j)-K(i,j)\right)^{2},$ (8)
    where $I$ is the reference image, $K$ is the reconstructed image, and $m$ and $n$ are the dimensions of the images. A higher value of PSNR indicate better performance.
  6. 6.
    Structural Similarity Index (SSIM): [51] The SSIM index is a method for predicting the perceived quality of digital television and cinematic pictures, as well as other kinds of digital images and videos. SSIM is designed to improve on traditional methods like PSNR and MSE, which have proven to be inconsistent with human eye perception. The SSIM index between two images $x$ and $y$ is defined as:
    $\text{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (9)
    where $\mu_{x}$ is the average of $x$, $\mu_{y}$ is the average of $y$, $\sigma_{x}^{2}$ is the variance of $x$, $\sigma_{y}^{2}$ is the variance of $y$, $\sigma_{xy}$ is the covariance of $x$ and $y$, and $C_{1}$ and $C_{2}$ are constants to stabilize the division with weak denominator. These last three metrics do not need the 3D ground truth and are widely used in literature [52, 53] for evaluation. SSIM ranges from -1 to 1, with higher values indicating better performance.
Precision-Recall curves:
Precision-recall curves are utilized to methodically evaluate how distance threshold $d$ changes influence precision $P(d)$ and recall $R(d)$ metrics, demonstrating the trade-off between these measurements under varying threshold conditions. To set the value of $d$ for the final assessment, we opt for a conservative estimate before the plateauing of precision-recall curves. For indoor scenarios, assuming a hypothetical grid size of 128x128x128 for reference, we establish $d$ at 0.005. In this scenario, the voxel size is calculated as $1/128\approx 0.0078125$, which makes the threshold of 0.005 smaller than the voxel size. This indicates a requirement for points to be closer than the dimensions of a single voxel to be identified as distinct, highlighting a prioritization of detail sensitivity within a hypothetically coarser grid. Such a setting is especially pertinent for capturing the complex geometries of indoor plants, where precision in detail is crucial. Due to the size and complexity of the scene, a threshold of 0.01 is selected for outdoor plant reconstructions.
2.5   Early Stopping of NeRF Training using LPIPS
In training NeRFs for plant scene reconstruction, the F1 score is essential for validating the accuracy of the reconstructed point cloud against the ground truth. The inherent challenge during the training phase of NeRFs is the absence of ground truth, paradoxically the output we aim to correspond. Moreover, the training process for NeRFs is notoriously compute-intensive. The cumulative costs become challenging when scaled to multiple scenes or across extensive agricultural fields.
Figure 4 shows the scatter plots of PSNR, SSIM, and LPIPS scores against the F1 score, alongside their respective Pearson correlation coefficients. This visualization offers an immediate visual assessment of the relationships between these metrics, and allows for a nuanced understanding of how accurately each metric predicts the true F1 score. The exceptionally strong negative correlation between LPIPS and F1 score (-0.82) reinforces the notion that LPIPS effectively captures the perceptual similarity between the reconstructed and ground truth point clouds, making it a reliable proxy for F1 score, the ultimate measure of reconstruction fidelity.
F1 Score
@@ 2.5   Early Stopping of NeRF Training using LPIPS
Figure 4: Correlation analysis between different metrics with F1 Score via Pearson coefficients: (a) PSNR, (b) SSIM, and (c) LPIPS.
  Scenario I    Scenario II    Scenario III
Algorithm 1 Plateau Detection Algorithm with LPIPS
1:procedure DetectPlateau($\mathcal{S},\mathcal{G},\theta,C$) $\triangleright$ Inputs: Sets of images per iteration $\mathcal{S}$, Sets of GT images $\mathcal{G}$, threshold $\theta$, consistency length $C$
2:     Initialize an empty list $\mathcal{M}$ to store average LPIPS values over training iterations.
3:     for each set $\mathcal{I}$ and corresponding GT set $\mathcal{G}_{I}$ in $\mathcal{S}$ and $\mathcal{G}$ do
4:         Initialize $\text{sumLPIPS}=0$.
5:         for each image $I$ and corresponding GT image $G$ in $\mathcal{I}$ and $\mathcal{G}_{I}$ do
6:              sumLPIPS += $\text{LPIPS}(I,G)$.
7:         end for
8:         Compute average LPIPS for the set: $\text{avgLPIPS}=\text{sumLPIPS}/|\mathcal{I}|$.
9:         Append avgLPIPS to $\mathcal{M}$.
10:     end for
11:     if $|\mathcal{M}|<C$ then
12:         return 00 $\triangleright$ Insufficient data for plateau detection
13:     end if
14:     for $i=1$ to $|\mathcal{M}|-1$ do
[…]
16:         for $j=\max(0,i-C+1)$ to $i$ do
17:              if $|\mathcal{M}[j]-\mathcal{M}[j-1]|\geq\theta$ then
[…]
19:              end if
20:         end for
21:         if consistent then
22:              return $i$triangleright Plateau point detected, Output: $P$
23:         end if
24:     end for
25:     return $|\mathcal{M}|-1$triangleright No plateau detected, Output: $P$
26:end procedure
The significant negative correlation between LPIPS and the F1 score (-0.82), PSNR (-0.81), and SSIM (-0.69) underscore the impact of LPIPS on the quality of 3D reconstruction (see supplementary material for detailed correlation matrix). The high magnitude of these coefficients, particularly the -0.82 with the F1 score, indicates that LPIPS is a robust predictor of reconstruction accuracy: as the perceptual similarity measure improves (meaning LPIPS decreases), the fidelity of the reconstructed point cloud to the ground truth improves correspondingly. This observation not only suggests the utility of LPIPS as a stand-in metric when the ground truth is unavailable but also highlights its potential as a more influential factor than traditional metrics such as PSNR (0.58) and SSIM (0.37) in determining the overall quality of NeRF-generated reconstructions.
Given this strong correlation, LPIPS emerges as a promising surrogate metric for early stopping during NeRF training. By monitoring LPIPS, one can infer the likely F1 score and make informed decisions about halting the training process. This method could decrease computational costs and time, as one need not await the completion of full training to predict its efficacy in terms of F1 score.
Algorithm for Plateau Detection: The plateau detection algorithm identifies a stabilization point in a series of metric values, such as LPIPS. The updated algorithm computes the average Learned Perceptual Image Patch Similarity (LPIPS) for each set of images in $\mathcal{S}$ against their corresponding ground truth images in $\mathcal{G}$. It then assesses the sequence of these average LPIPS values to identify a plateau, using a specified threshold $\theta$ and a consistency length $C$. The detection of the plateau point $P$ is crucial for indicating an optimal stopping point in the training process. To validate the efficacy of the early stopping algorithm, we applied it to a diverse dataset comprising five plant types captured in both indoor and outdoor settings. The threshold ($\theta$) was set to 0.005, and the consistency length (C) was fixed at 6. The granularity of interpolation was set to 1000, spanning a total of 60000 training iterations. These hyperparameters were chosen based on empirical observations to ensure a balance between computational efficiency and reconstruction accuracy.
3   Results
We evaluated the performance of NeRF models across various scenarios, from controlled indoor environments to complex outdoor field conditions, using key performance metrics to assess their efficacy in 3D plant reconstruction. The NeRFs were trained on an NVIDIA A100 GPU with 80GB GPU RAM attached to an AMD EPYC 7543 32-core CPU with 503GB CPU RAM. Post-training, the models are converted into point clouds with approximately a million points each. Estimated camera poses from COLMAP are visualized in Figure 5, and a summary of the performance metrics of each of the three scenarios is given in Table 2. 3D evaluation metrics are presented in this section; for a more granular analysis of 2D image metrics, please refer to Supplement. Visually, the performance of each model could be assessed using Precision and Recall as shown in Figure 6. The Precision-Recall curves of the different Scenarios for different threshold values are shown in Figure 7.
Visualization Color Code: The color-coded visualizations employed provide an intuitive understanding of spatial relationships within the 3D reconstructed plant structures. The interpretation of colors is as follows:
  • Grey: (Correct) Represents points within a predefined distance threshold relative to the reference point cloud. This color indicates accurate points in precision and recall evaluations, where precision assesses the reconstruction against the ground truth, and recall evaluates the ground truth against the reconstruction.
  • Red: (Missing) Depicts points in the point cloud being tested that are beyond the distance threshold but within 3 standard deviations from the nearest point in the reference point cloud. These points are considered inaccuracies, showing missing details in the reconstruction when assessing precision and highlighting missing elements in the ground truth during recall analysis.
  • Black: (Outlier) Highlights points in the point cloud being tested that are more than 3 standard deviations away from any point in the reference point cloud. These points are extreme outliers and represent significant errors in the reconstruction relative to the ground truth for precision evaluations, and similarly significant discrepancies in the ground truth relative to the reconstruction for recall.
(a) Scenario-I.
(b) Scenario-II.
(c) Scenario-III.
Figure 5: Camera pose estimations across three different scenarios. (a) Scenario I, (b) Scenario II, (c) Scenario III.
Table 2: Performance metrics of NeRFs reconstruction techniques across scenarios I, II and III.
# Model Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s) $\downarrow$
Instant-NGP 24.66 90.62 38.77 23.41 0.81 0.17 756
I TensoRF 9.58 43.34 15.69 14.69 0.55 0.66 1973
NeRFacto 73.57 94.72 82.81 22.24 0.73 0.12 1938
Instant-NGP 23.45 58.57 33.49 19.08 0.64 0.31 1886
II TensoRF 20.5 55.34 29.91 15.54 0.42 0.56 2607
NeRFacto 64.47 76.8 70.1 18.93 0.64 0.25 1226
Instant-NGP 15.06 59.55 24.04 18.54 0.47 0.4 1466
III TensoRF 40.95 75.62 53.13 17.32 0.39 0.55 1965
NeRFacto 68.29 82.32 74.65 16.7 0.32 0.34 1499
3.1   Scenario I - Single Plants Indoors
We first look at the results of reconstructing a single plant in an indoor environment. Detailed evolution of each metric over training iterations is given in the Supplement.
Precision: For Scenario I, NeRFacto, achieved the highest precision followed by TensoRF and Instant-NGP (see Figure 6) after 30000 iterations. Across all models, precision generally increases with the number of iterations. For a detailed evaluation of the change of precision with iterations, please refer the Supplement.
{tblr}
colspec = X[0.015,c,t]X[0.29,c,t]X[0.29,c,t]X[0.29,c,t], stretch = 0, rowsep = 3pt, Precision & Refer to caption Refer to caption Refer to caption
Recall Refer to caption Refer to caption Refer to caption
Scenario I
Precision Refer to caption Refer to caption Refer to caption
Recall Refer to caption Refer to caption Refer to caption
Scenario II
Precision Refer to caption Refer to caption Refer to caption
Recall Refer to caption Refer to caption Refer to caption
Scenario III
Instant-NGP TensoRF NeRFacto
[…]
Recall: The recall metric follows a similar trend, with Instant-NGP and NeRFacto showing increases with more iterations, indicating an enhanced ability to encompass points from the ground truth. Notably, NeRFacto achieves remarkably high recall values (over 90) at higher iterations, suggesting its superiority in the completeness of reconstruction. TensoRF’s recall values are significantly lower, indicating that it may miss more details from the ground truth compared to the other models.
F1 Score: The F1 score, balancing precision and recall, highlights NeRFacto as the most balanced model, especially at higher iterations, with scores above 80. Instant-NGP shows a significant improvement in F1 scores as iterations increase, but it doesn’t reach the same peak as NeRFacto. TensoRF lags in this metric, indicating a less balanced performance between precision and recall.
Computation Time: Time efficiency is a crucial factor, especially for practical applications. Instant-NGP demonstrates a relatively balanced approach between efficiency and performance, with time increments correlating reasonably with the increase in iterations. However, it becomes time-consuming at high iterations (20000 and 30000). NeRFacto, while showing better performance in many metrics, demands considerably more time, especially at higher iterations, which could be a limiting factor in time-sensitive scenarios. The evolution of precision over training time for NeRFacto is given in the supplementary material. TensoRF, despite its lower performance in other metrics, maintains a more consistent time efficiency, suggesting its suitability for applications where time is a critical constraint. Please see the Supplement for the evolution of precision metric over the training iterations.
Overall Performance and Suitability: In sum, NeRFacto emerges as the most robust model in terms of precision, recall, F1 score, and image quality metrics (PSNR, SSIM, LPIPS), making it highly suitable for applications demanding high accuracy and completeness in 3D modeling. However, its time inefficiency at higher iterations might restrict its use in time-sensitive contexts. Instant-NGP presents a good balance between performance and efficiency, making it a viable option for moderately demanding scenarios. Detailed results are given in Table 2, after complete training. For more granular look of each metric value over the training iteration for all the algorithms, consult the supplementary. The Precision-Recall curves based on varying distance threshold after maximum training of 30,000 iterations is given in Figure 7.
Scenario I
# of points (%)
Threshold (m)
Threshold (m)
Threshold (m)
Scenario II
# of points (%)
Threshold (m)
Threshold (m)
Threshold (m)
Scenario III
# of points (%)
Threshold (m)
@@ 3.1   Scenario I - Single Plants Indoors
Figure 7: Precision-Recall curves for the three scenarios based on varying distance threshold after 30,000 iterations (Scenario I and II) and 60,000 iterations (Scenario III).
@@
Insight 1: Computational Cost and Accuracy Trade-off in Instant-NGP and NeRFacto: The steep increase in performance metrics with the number of iterations for both Instant-NGP and NeRFacto suggests that these models require a substantial amount of data processing to achieve high accuracy, which is critical in high-fidelity 3D modeling. However, this also implies a higher computational cost, which needs to be considered in practical applications.
@@
Insight 2: Model Suitability in High-Detail 3D Reconstructions: The significant disparity in the performance of TensoRF compared to the other two models, particularly in precision and recall, indicates that not all NeRF models are equally suited for tasks requiring high-detail 3D reconstructions. This highlights the importance of model selection based on the specific requirements of the application.
@@
Insight 3: Divergence in 2D Image Quality and 3D Reconstruction in Instant-NGP: A detailed examination reveals that Instant-NGP demonstrates strength in 2D image quality metrics such as PSNR, SSIM, and LPIPS, reflecting its ability to produce better rendered image quality. However, this excellence in 2D imaging does not correspondingly extend to 3D reconstruction metrics like Precision, Recall, and F1 Score. This observation highlights a significant distinction in the challenges associated with optimizing for high-quality image rendering as opposed to achieving accurate 3D representations. The model’s adeptness at rendering highly detailed 2D images does not necessarily imply its effectiveness in accurately reconstructing complex 3D structures, particularly in the context of intricate plant models. This insight underscores the need for a nuanced approach in evaluating the performance of models that are tasked with both 2D image rendering and 3D spatial reconstruction.
3.2   Scenario II - Multiple Plants Indoors